# Learned Contextual LiDAR Informed Visual Search in Unseen Environments

### v1

@@ Abstract
This paper presents LIVES: LiDAR Informed Visual Search, an autonomous planner for unknown environments. We consider the pixel-wise environment perception problem where one is given 2D range data from LiDAR scans and must label points contextually as map or non-map in the surroundings for visual planning. LIVES classifies incoming 2D scans from the wide Field of View (FoV) LiDAR in unseen environments without prior map information. The map-generalizable classifier is trained from expert data collected using a simple cart platform equipped with a map-based classifier in real environments. A visual planner takes contextual data from scans and uses this information to plan viewpoints more likely to yield detection of the search target. While conventional frontier based methods for LiDAR and multi sensor exploration effectively map environments, they are not tailored to search for people indoors, which we investigate in this paper. LIVES is baselined against several existing exploration methods in simulation to verify its performance. Finally, it is validated in real-world experiments with a Spot robot in a 20x30m indoor apartment setting. Videos of experimental validation can be found on our project website at https://sites.google.com/view/lives-icra-2024/home.
@@ I INTRODUCTION
Autonomous planning and real-world execution for active search and exploration are receiving significant attention from the robotics community due to their relevance in several scenarios including inspection, surveillance, and Search and Rescue (SAR) [27]. Autonomous ground robots cannot assume reliable map information a priori during active search missions, for example in a disaster response scenario. This demands environment perception tailored for the task of locating any persons in need quickly. Today’s robots are frequently equipped with LiDAR sensors that cast a full view of the surroundings. However, use of these scans is often limited to the task of localization and mapping. This paper addresses the exploitation of wide Field of View (FoV) LiDAR scans for visual planning. We propose to classify LiDAR scan points as map or non-map in the robot’s surroundings for improved visual search planning.
@@
Several methods exist for pixel-wise classification of 3D LiDAR scans [22, 23, 34]. However, [15] notes that these methods are limited by the high computational requirements to handle to large input size of 3D cloud data. In contrast, real time projection-based methods aim to detect moving obstacles [13] or objects to be avoided in 3D scans, often for autonomous driving applications [18, 5, 20]. Despite parallel improvements in efficiency of planning algorithms for both LiDAR and vision sensors [6, 14, 28, 32] it remains an open problem to enhance performance in autonomous planning and execution for the task of visual search. There exists a gap of real-time pixel-wise classification of LiDAR scans for active visual search tasks. This work addresses this limitation with an online 2D pixel-wise LiDAR scan classifier and context informed visual planner. In particular, classified LiDAR scans are used to generate highly informative plans quickly with a multisensor Frontier Exploration [33] based planner.
@@
Figure 1: Overview of LIVES. In (a) raw 2D LiDAR scans, pose information and image data are received from the Spot robot. They are sent through a neural network classifier III-C (b) to label each point as map or non-map (c). Separately, LiDAR scans and pose data are used to maintain the global search maps (d). Global search maps are used to generate frontier points (e), which are used for finding candidate viewpoints. Finally, each frontier point is sampled at four viewpoints (f) and scored III-E2. In this toy example the left viewpoint in (f) will be scored highest and selected as the next waypoint due to the high number of unknown cells and non-map points observed. Finally, the selected viewpoint is sent back to the robot for execution.
Active planners in unknown spaces are required to make decisions based on incomplete and noisy information about partial environments. State of the art visual task planners like [6, 32] account for LiDAR scans without considering contextual information. We provide a visual planner with map and non-map data in the surroundings found in wide FoV LiDAR scans. This is accomplished by learning to distinguish points unlikely to be permanent features in unknown environments. The visual planner gives priority to features of interest under the assumption that the search target belongs to the set of non-map points. This information obviates the need to visually inspect parts of the environment that are likely to be permanent. An overview of the proposed method can be found in Fig. 1.
@@
  • A learning strategy for pixel-wise scan classification as map and non-map in unseen, real-world environments
  • A formulation that incorporates classified scan information into a planner for visual search
@@
  • Deploy and verify the proposed method on a real robotic system in an unseen indoor environment
Sections II, III and IV describe related works, methods an results including baseline comparisons.
@@ II-A Lidar Classification
Indoor pointcloud classification is tackled in 3D [30] using the KPConv architecture [29]. This method requires the offline annotation and map building at instantiation, and is used for localization and navigation in later work [31]. Neither work validates on a real robot platform. Ref. [21] presents an in depth comparison and framework for Iterative Closest Point (ICP) methods, similar to that of [30], focused on map building. Finally, [26] present a map-free framework for segmenting 3D pointclouds into ground, edges and planar points, which are used in a feature extraction process for mapping environments. In contrast LIVES recognizes non-map features.
II-B Next-Best View and Informative Path Planning
Exploration is often solved using next-best viewpoint planners like Frontier Exploration [33, 14]. A common approach uses an active perception formulation to maximize entropy reduction [17, 12]. Ref. [11] learns a policy to generate non-myopic plans online by estimating potential gains of exploring different areas. Ref. [19] similarly learn an information aware policy offline that proposes waypoints for the robot. A local planner then generates dynamically feasible trajectories based on these suggestions. In contrast to the proposed approach, [11, 19] only provide simulation results. Similar to our method, [36] focuses on the problem of visual search. They propose an end-to-end policy that learns to generate discrete actions that lead a robot to a prescribed viewpoint in the given environment, however, the action space is limited to simple discrete actions and they do not leverage LiDAR information. Ref. [35] propose a POMDP planner for generalized local object search in cluttered 3D environments and extend it with a 2D planner for larger areas. In contrast to LIVES, this work is aimed at small areas and small objects, rather than quickly searching buildings for people.
@@ II-C World Prediction Based Planning
Often world prediction methods involve learning a model for relevant environments and leveraging their structure for exploring high-value regions first. Ref. [25] use topological features from a database of worlds to inform a frontier-based exploration policy and increase the rate of area explored. Similarly, [28] uses a prediction module that detects semantic objects, classifies frontiers, and predicts information gain beyond those frontiers. A planner uses this to explore rooms fully before continuing global exploration, unlike LIVES that avoids complete exploration to reduce search completion time. Ref. [24] instead consider visual navigation with occupancy anticipation from egocentric RGB-D based visual context. Ref. [10] instead learn to predict locations of exits in building with a convolutional neural network (CNN) using a database of building blueprints. Prediction guide the robot towards likely exit locations.
@@ III-B2 STF
When Eq. 2 is greater than some threshold, point $p_{i}^{j}$ is classified as an STF. Remaining points in $p_{i}^{j}$ are classified as DFs. The result is classified scan $s^{\text{class}}_{i}$. In practice, there are a large number of STFs and DFs due to the size of scans, therefore it is important to filter them. First, the pooling operator is used to reduce duplicates within a radius. Second, pooled STFs and DFs within some distance of LTFs are removed to eliminate false positives caused by localization drift. Finally, points inside visually observed regions of the map are removed.
@@ III-C Map-free LiDAR Scan Classification
The classification method defined in III-B is limited by the necessity of static map information $\mathcal{M}$. Therefore, this work uses supervised learning to train a model to reproduce classifications for $\{p^{j}_{i}\}_{j=1:n_{i}}$ in scan $s_{i}$. Similarly to the ground-truth, non-map points are filtered by pooling and then by removing those in previously inspected regions.
III-C1 Dataset and Data Acquisition
The collected dataset $\mathcal{D}$ consists of tuples of data from the ground truth classification model. The tuples are given as $\mathcal{D}=\{(x_{i},s_{i},s_{i}^{\text{class}})\}^{N}_{i=1}$, where $x_{i}$ represents robot pose, $s_{i}$ is the raw LiDAR scan data, and $s_{i}^{\text{class}}$ is the classified scan. $N$ is the total number of data points collected. Collecting a dataset using a real robot requires a human expert to control the robot over long periods of time. This approach is difficult to scale due to mechanical limits of the robot and difficulty of control over large, potentially out-of-sight places. To ease data acquisition we design a steerable cart platform (see fig. 3) to mimic the robot. Human operators easily maneuver this cart to generate the dataset without strain to the robot. We mount a Velodyne Puck LiDAR for recording scan information and a front-mounted Intel RealSense T265 stereo camera to obtain the cart’s odometry estimate, required for running the ground-truth classification. Finally, a laptop is used to maintain the pose estimate [9], collecting the data and generating online ground-truth labels for the dataset. This laptop records 1) robot pose at each timestep, 2) raw scan data, and 3) classified scan from the ground-truth module for dataset $\mathcal{D}$. The Puck has $n_{i}=897$ and classified scans are recorded at 5Hz. A dataset of size $N\simeq 145,000$ is collected in this study, representing roughly 8 hours of time. In practice, the model only receives raw scan data, pose estimate and previously predicted labels from the robot during operation and must classify pixel-wise the scan online.
Figure 3: A figure depicting the cart used for labeled data acquisition and the Spot robot used for deployment of the proposed method. The RealSense provides odometry estimate to the cart platform, required for localization and ground-truth estimation. The spot is equipped with an RGB-D Azure Kinect for detection during the search task.
III-C2 Architecture and Training
We formulate pixel-wise LiDAR classification as a supervised learning problem during training and an auto-regressive problem during inference. Given dataset $\mathcal{D}$, we shuffle and batch the dataset between time steps and locations to minimize location-based and temporal bias during training. Because no map information is provided explicitly, we provide the model with a history buffer of $k$ time steps for scan ranges and robot poses, and $k-1$ previously estimated ground truth labels. Because we do not have access to ground truth labels at inference time, we corrupt the ground truth labels during training by randomly bit-flipping the classification of 10% of the labels, chosen by sampling indices from a uniform random distribution. We then take the exponential weighted average of the last $k-1$ corrupted pixel-wise classifications to estimate the current pixel-wise classification. This estimate is concatenated with the $k-1$ corrupted ground truth labels to produce a $k$-length matrix of estimated scan labels. During inference, the label history buffer is populated by previous estimates from the policy, cropped by a length $k$ sliding-window. The history buffer is then concatenated with the updated pixel-wise exponential weighted average. At the initialization of inference, the policy is bootstrapped with zero-value poses, scans, and labels, and run for $k$ steps of inference until the history buffer is initialized.
The policy consists of a temporal-convolutional encoder (TCN) for each input. Each encoder contains a single convolutional layer with scan-wise circular padding and a single linear layer. The scan and label encoders have kernel size $[k,k]$, and the pose encoder has kernel size $[1,3]$. Hyperbolic tangent activations are used for all layers. The output of each encoder is the same size as the LiDAR scan to be classified. The output of the pose encoder is summed with the output of the scan encoder as a pose correction, and then normalized by applying a hyperbolic tangent function. The resulting LiDAR scan encoding is multiplied element-wise with the label encoding. A threshold is applied to the raw logits such that each pixel is classified as either a map point or a non-map point. The policy is trained with mean squared error loss against ground truth classifications. Training for 20 epochs on a workstation Nvidia 3080 12Gb takes 3 minutes. The full model architecture is shown in Fig. 2.
@@ III-E1 Candidate Viewpoint Generation
Frontier points in each of the two global search maps are generated using a process similar to [33], shown in Fig. 1(e). Any free cell ($e=0$) adjacent to an unknown cell ($e=-1$) is a frontier edge, which are grouped together into frontiers. Candidate viewpoints at frontiers are similar to those of [32]. First, a set of vision based viewpoints are generated by clustering frontiers on the visual global search map. Next, a set of LiDAR based viewpoints are generated by clustering frontiers on the LiDAR global search map. Each frontier cluster becomes four candidate viewpoints, facing each direction, to account for the limited FoV of the vision sensor.
@@ IV Results
We test three hypotheses through our evaluation: (1) The proposed approach to inform an exploring agent with non-map features is effective at finding search targets, (2) LIVES reduces detection time compared against existing exploration algorithms, and (3) The proposed method can be deployed on real robot hardware and used in a human environment.
@@ IV-B Simulation and Baseline Comparison
Simulations are performed using Gazebo-ROS on the Turtlebot [4]. ROS move base is used for navigation [1]. Baselines are: 1) a lightly modified version of [32] (Multisensor Frontier Exploration), 2) the Next-Best View Planner (NBVP) [7], 3) a Rapidly-Exploring Random Tree (RRT) [3], and 4) LIVES given ground-truth scan classifications. The Multisensor baseline is equivalent to LIVES and ground-truth without contextual LiDAR information.
@@
Two Gazebo world environments are tested with easy and hard setups in each. The first map is the apartment (20x30m) and examines the impact of non-target objects on LIVES (see Fig. 5(a-b)). Here the easy setting incorporates fewer objects in the environment and the search target being closer to the robot while hard is defined by more objects in the environment and larger distance to the target. The second map is the office (25x45m) and tests the method in a large environment under significantly varied target position (see Fig. 5(c)). The easy and hard settings are defined by the search target being in the same or different half of the environment as the robot starting position. Both difficulty levels in each map are repeated 10 times. Average detection times and success rates are reported in Table I. The time limit given to planners is 2 and 3 minutes in the apartment easy and hard settings then 3 and 5 minutes in the office easy and hard settings. Trials where the planner fails or the time limit is reached are reported as a failure and failed trials are reported as the maximum allotted time.
@@
Results indicate that LIVES, the Ground-Truth variant and the Multisensor Frontier Exploration method significantly outperform the RRT and NBVP baselines in all settings except for the hard office setting, where the RRT method performs similarly to LIVES and Ground-Truth methods. In the easy apartment setting we observe similar performance among the three planners. However in the hard setting, where LIVES and Ground-Truth perform nearly identically, there is roughly a 15% improvement with the addition of the knowledge of non-map points compared with the Multisensor method. We see a large gap, however, in the easy office setting with a near 30% improvement with the addition of contextual LiDAR information over the Multisensor variant. In the hard office setting, LIVES and Ground-Truth variant outperform the RRT method by 5-12%. Overall, the results indicate that the proposed method to inform the frontier planner with non-map information outperforms the baselines across the full spectrum of environments. We hypothesize this is due to the requirement for the exploration planners to exhaustively explore the environment, whereas the proposed method focuses on non-map points, allowing the agent to move on when the target is unlikely to be nearby. These results support hypotheses (1) and (2) that contextual LiDAR information significantly improves performance in the search task.
@@
Figure 6: A time lapse of one experiment scenario. In (a) the robot pose is a purple arrow with the visual costmap, global search map, frontier points and non-map points shown. The raw scan data is shown as small white squares. The target object is behind the robot starting position. Figure (b) occurs immediately before a planning step, wherein the marked frontier is selected due to the high occurrence of non-map points nearby. (c) occurs after the planning step and results in detection.

### v3

@@ Abstract
This paper presents LIVES: LiDAR Informed Visual Search, an autonomous planner for target search in unknown environments. We consider the pixel-wise environment perception problem where one is given wide Field of View 2D scan data and must perform LiDAR segmentation to contextually label points in the surroundings. These pixel classifications provide an informed prior on which to plan next best viewpoints during visual search tasks. The map-generalizable classifier is trained from expert data collected using a simple cart platform equipped with a map-based classifier. An autonomous exploration planner takes the contextual data from scans and uses that prior to plan viewpoints more likely to yield detection of the search target. In order to achieve this, we propose a utility function that accounts for traditional metrics like information gain and path cost and also for the additional contextual information from the scan classifier. LIVES is baselined against several existing exploration methods in simulation to verify its performance. Finally, it is validated in real-world experiments searching for single and multiple targets with a Spot robot in two unseen environments. Videos of experimental validation, implementation details and open source code can be found on our project website at https://sites.google.com/view/lives-2024/home.
@@ I INTRODUCTION
Planning and real-world execution for autonomous visual exploration are receiving significant attention from the robotics community due to their relevance in a wide range of scenarios including inspection [35], target search [1, 12], disaster response [15, 40, 4], agricultural monitoring [13], reconnaissance and surveillance [16, 20, 23]. Active planners in unknown spaces are required to make decisions based on incomplete and noisy information about partial environments. This requires planners equipped to generate high quality plans under uncertainty. Conventional methods for LiDAR and multi sensor exploration effectively map environments with focus on completeness, rather than being tailored time critical missions, which we investigate in this paper.
@@
Today’s robots are frequently equipped with LiDAR sensors that cast a full view of the surroundings. Use of scans is often limited to obstacle avoidance, localization and mapping. In contrast, state of the art exploration planners like [3, 42] demonstrate clear improvements by fusing LiDAR and visual information. At the same time, the potential to use additional information from LiDAR scans for online control has grown with access to parallel computing and more sophisticated GPUs. While several methods exist for 3D LiDAR scan segmentation [29, 30, 45], ref. [17] notes these methods are limited by the high computational requirements to handle to large input size of 3D cloud data. In contrast, real time projection-based methods aim to detect moving obstacles [11] or objects to be avoided in 3D scans, often for autonomous driving applications [21, 2, 26]. There are also numerous planners designed to reduce uncertainty in new environments by exploiting features common to the environments in which they are designed to operate [33, 31]. Despite improvements in efficiency of planning algorithms for both LiDAR and vision sensors [3, 14, 36, 42], it remains an open problem to enhance performance in autonomous planning and execution for the task of visual search [43]. There exists a gap of visual search planners utilizing additional information available in LiDAR scans from segmentation. This work combines these concepts to reduce uncertainty in indoor environments by segmenting 2D LiDAR scans online for utilization by a next-best view (NBV) planner.
@@
Figure 1: Overview of a planning step of LIVES. In (a) raw 2D LiDAR scans, pose information and image data are received from the Spot robot. They are sent through a neural network classifier III-C (b) to label each point as map or non-map (c). Separately, LiDAR scans and pose data are used to maintain the global search maps (d). Global search maps are used to generate frontier points (e), which are used for finding candidate viewpoints. Finally, each frontier point is sampled at four viewpoints (f) and scored III-E2. In this toy example the left viewpoint in (f) will be scored highest and selected as the next waypoint due to the high number of unknown cells and non-map points observed. Finally, the selected viewpoint is sent back to the robot for execution.
The key insight in this work is to exploit contextual information available in wide Field of View (FoV) LiDAR scans to filter out points unlikely to be the search target. In particular, this method deviates from traditional exploration methods, which focus on completeness. Rather, LIVES prioritizes regions more likely to be the search target in the robot’s surroundings. We hypothesize exploring agents can improve time efficiency in visual search by exploiting this additional information. We accomplish this using a map based classifier to train a model to recognize features inherent to many structured, indoor environments. The perception model feeds the planner those points unlikely to be permanent, non-map points, for use during viewpoint planning. These points guide the NBV planner towards possible search target locations, assumed to be in the set of non-map points in the environment. Priority is given to features of interest using a novel utility function that accounts for both traditional metrics and non-map points. This information obviates the need to visually inspect parts of the environment that are likely to be permanent. An overview of the proposed method can be found in Fig. 1.
@@
  • A learning strategy for online scan segmentation as map and non-map in unseen, real-world environments
  • A formulation that incorporates classified scan pixel information into a planner for visual search
@@
  • Deploy and verify the proposed method on a real robotic system in multiple unseen indoor environments
@@ II-A Lidar Classification
Indoor pointcloud classification is tackled in 3D [38] using the KPConv architecture [37]. This method requires the offline annotation and map building at instantiation, and is used for localization and navigation in later work [39]. Neither work validates on a real robot platform. The proposed work instead focus on low computation 2D scan pixel classification.Ref. [28] presents an in depth comparison and framework for Iterative Closest Point (ICP) methods, similar to that of [38], focused on map building rather than contextual labeling for planning. Finally, [34] present a map-free framework for segmenting 3D pointclouds into ground, edges and planar points, which are used in a feature extraction process for mapping outdoor environments. In contrast LIVES recognizes non-map features for indoor environments.
II-B Next-Best View and Multisensor Planning
Exploration is often solved using next-best viewpoint planners like Frontier Exploration [44, 14]. A common approach uses an active perception formulation to maximize entropy reduction [19, 10]. Ref. [12] combine frontier and sampling based methods to reduce computational overheard for deployment on limited compute UAVs. Ref. [9] learns a policy to generate non-myopic plans online by estimating potential gains of exploring different areas. Ref. [22] similarly learn an information aware policy offline that proposes waypoints for the robot. These methods demonstrate efficient exploration in simulation and on hardware, however none are explicitly designed for visual search using contextual information. Ref. [46] propose a POMDP planner for generalized local object search in cluttered 3D environments and extend it with a 2D planner for larger areas. In contrast to LIVES, their work is aimed at 3D search for small objects, rather than quickly searching buildings for people.
Multisensor exploration involves the inclusion of LiDAR information in planning for visual exploration tasks. State of the art multisensor exploration with multiple UAVs [3] leverages the single agent volumetric exploration algorithm presented in [42]. These methods combine frontier generation for both sensors and use a single utility function to score frontier candidates for surface mapping. Ref. [27] maintains a dense 3D pointcloud for recognizing of objects in subterranean tunnel environments. These classifications are utilize by a Pan-Tilt-Zoom camera to perform object detection in Search and Rescue applications. These multisensor approaches demonstrate efficient results in a number of real world environments, however none of these methods is tailored for the indoor environments considered in this article.
@@ II-C World Prediction Based Planning
Often world prediction methods involve learning a model for relevant environments and leveraging their structure for exploring high-value regions first. Ref. [33] use topological features from a database of worlds to inform a frontier-based exploration policy and increase the rate of area explored. Prediction involves recognizing features common to tunnel networks to improve exploration strategies. Similarly, [36] uses a prediction module that detects semantic objects, classifies frontiers, and predicts information gain beyond those frontiers. The planner uses this to explore rooms fully before continuing global exploration. This contrasts from the proposed method that aims to ignore low value regions in order to reduce search completion time. Ref. [31] instead consider visual navigation with occupancy anticipation from egocentric RGB-D based visual context. Ref. [8] instead learn to predict locations of exits in building with a convolutional neural network (CNN) using a database of building blueprints. Prediction guide the robot towards likely exit locations. This article describe a novel method for world prediction from 2D scans.
@@ III-B2 STF
When Eq. 2 is greater than some threshold, point $p_{i}^{j}$ is classified as an STF. Remaining points in $p_{i}^{j}$ are classified as DFs. The result is classified scan $s^{\text{class}}_{i}$.
@@ III-C Map-free LiDAR Scan Classification
The classification method defined in III-B is limited by the necessity of static map information $\mathcal{M}$. Therefore, this work uses supervised learning to train a model to reproduce classifications for $\{p^{j}_{i}\}_{j=1:n_{i}}$ in scan $s_{i}$.
III-C1 Dataset and Data Acquisition
The collected dataset $\mathcal{D}$ consists of tuples of data from the ground truth classification model. The tuples are given as $\mathcal{D}=\{(x_{i},s_{i},s_{i}^{\text{class}})\}^{N}_{i=1}$, where $x_{i}$ represents robot pose, $s_{i}$ is the raw LiDAR scan data, and $s_{i}^{\text{class}}$ is the classified scan. $N$ is the total number of data points collected. Collecting a dataset using a real robot requires a human expert to control the robot over long periods of time. This approach is difficult to scale due to mechanical limits of the robot and difficulty of control over large, potentially out-of-sight places. To ease data acquisition we design a steerable cart platform (see website for details) to mimic the robot. Human operators easily maneuver this cart to generate the dataset without strain to the robot. We mount a Velodyne Puck LiDAR for recording scan information and a front-mounted Intel RealSense T265 stereo camera to obtain the cart’s odometry estimate, required for running the ground-truth classification. Finally, a laptop is used to maintain the pose estimate [7], collecting the data and generating online ground-truth labels for the dataset. This laptop records 1) robot pose at each timestep, 2) raw scan data, and 3) classified scan from the ground-truth module for dataset $\mathcal{D}$. The Puck has $n_{i}=897$ and classified scans are recorded at 5Hz. A dataset of size $N\simeq 145,000$ is collected in this study, representing roughly 8 hours of time. In practice, the model only receives raw scan data, pose estimate and previously predicted labels from the robot during operation and must classify pixel-wise the scan online. Training data can be found on the project website.
III-C2 Architecture and Training
We formulate pixel-wise LiDAR classification as a supervised learning problem during training and an auto-regressive problem during inference. Given dataset $\mathcal{D}$, we shuffle and batch the dataset between time steps and locations to minimize location-based and temporal bias during training. Because no map information is provided explicitly, we provide the model with a history buffer of $k$ time steps for scan ranges and robot poses, and $k-1$ previously estimated ground truth labels. Because we do not have access to ground truth labels at inference time, we corrupt the ground truth labels during training by randomly bit-flipping the classification of 10% of the labels, chosen by sampling indices from a uniform random distribution. We then take the exponential weighted average of the last $k-1$ corrupted pixel-wise classifications to estimate the current pixel-wise classification. This estimate is concatenated with the $k-1$ corrupted ground truth labels to produce a $k$-length matrix of estimated scan labels. During inference, the label history buffer is populated by previous estimates from the policy, cropped by a length $k$ sliding-window. The history buffer is then concatenated with the updated pixel-wise exponential weighted average. At the initialization of inference, the policy is bootstrapped with zero-value poses, scans, and labels, and run for $k$ steps of inference until the history buffer is initialized. In practice there are a large number of non-map points due to the size of scans ($n_{i}=897$), therefore we apply filtering steps like pooling and removing those in previously inspected areas before these points are used for planning. Code for training a model can be found on the project website.
@@ III-C2 Architecture and Training
The full model architecture is shown in Fig. 2. The policy consists of a temporal-convolutional encoder (TCN) for each input. Each encoder contains a single convolutional layer with scan-wise circular padding and a single linear layer. The scan and label encoders have kernel size $[k,k]$, and the pose encoder has kernel size $[1,3]$. Hyperbolic tangent activations are used for all layers. The output of each encoder is the same size as the LiDAR scan to be classified. The output of the pose encoder is summed with the output of the scan encoder as a pose correction, and then normalized by applying a hyperbolic tangent function. The resulting LiDAR scan encoding is multiplied element-wise with the label encoding. A threshold (positive/negative) is applied to the raw logits such that each pixel is classified as either a map point or a non-map point. The policy is trained with mean squared error loss against ground truth classifications. Training for 20 epochs on a workstation Nvidia 3080 12Gb takes 3 minutes.
@@ III-E1 Candidate Viewpoint Generation
Frontier points in each of the two global search maps are generated using a process similar to [44], shown in Fig. 1(e). Any free cell ($e=0$) adjacent to an unknown cell ($e=-1$) is a frontier edge, which are grouped together into frontiers. Any group of frontier points beyond a minimum size becomes its own frontier. Frontier centroids of each frontier in both the visual and LiDAR search maps are found by clustering. Four candidate viewpoints are evaluated at each frontier centroid to account for the limited FoV of the vision sensor.
@@ IV Results
We test three hypotheses through our evaluation: (1) The proposed approach to inform an exploring agent with non-map features is effective at finding search targets, (2) LIVES reduces detection time compared against existing exploration algorithms, and (3) The proposed method can be deployed successfully on real robot hardware in unseen indoor environments.
@@ IV-B Simulation and Baseline Comparison
Simulations are performed using Gazebo-ROS on the Turtlebot [32]. ROS move base is used to navigate to the goal points given from the viewpoint selection methods compared [25]. Baselines are: 1) a lightly modified version of [42] (Multisensor Frontier Exploration), 2) the Next-Best View Planner (NBVP) [5], 3) a Rapidly-Exploring Random Tree (RRT) [41], and 4) LIVES given ground-truth scan classifications. Ref. [42] described a highly efficient planner for multisensor 3D exploration in underwater environments that has also been extended to multi-drone exploration [3]. The modifications from the original paper are as follows. First, the method is projected into 2D. Second, candidate viewpoint selection is changed. Originally, candidate viewpoints are selected surrounding frontier centroids, but in this work the candidate viewpoints are selected at centroids, with candidate views in four directions. This modified version of the multisensor frontier method is equivalent to the proposed approach if our method is not given any contextual information from the scan classifier.
@@
Two Gazebo world environments are tested with easy and hard setups in each. The first map is the apartment (20x30m) and examines the impact of non-target objects on LIVES (see Fig. 4(a-b)). Here the easy setting incorporates fewer objects in the environment while hard is defined by more objects in the environment. The second map is the office (25x45m) and tests the method in a large environment under significantly varied target position (see Fig. 4(c)). The easy and hard settings are defined by the search target being in the same or different half of the environment as the robot starting position. Both difficulty levels in each map are repeated 10 times. Average detection times and success rates are reported in Table I. The time limit given to planners is 2 and 3 minutes in the apartment easy and hard settings then 3 and 5 minutes in the office easy and hard settings. Trials where the planner fails or the time limit is reached are reported as a failure and failed trials are reported as the maximum allotted time.
@@
Results indicate that LIVES, the Ground-Truth variant and the Multisensor Frontier Exploration method significantly outperform the RRT and NBVP baselines in all settings except for the hard office setting, where the RRT method performs similarly to LIVES and Ground-Truth methods. In the easy apartment setting we observe similar performance among the three planners. However in the hard setting, where LIVES and Ground-Truth perform nearly identically, there is roughly a 15% improvement with the addition of the knowledge of non-map points compared with the Multisensor method. There exists a large gap in the easy office setting with a near 30% improvement with the addition of contextual LiDAR information over the Multisensor variant. In the hard office setting, LIVES and Ground-Truth variant outperform the RRT method by 5-12%. Overall, the results indicate that the proposed method to inform the NBV planner with non-map information outperforms the baselines across the full spectrum of environments. We hypothesize this is due to the requirement for the exploration planners to exhaustively explore the environment, whereas the proposed method focuses on non-map points, allowing the agent to move on when the target is unlikely to be nearby. These results support hypotheses (1) and (2) that contextual LiDAR information significantly improves performance in the search task.
Figure 5: Ground truth maps of the two experimental environments. (a) shows the 15mx38m hallway environment and (b) shows the 20mx30m apartment.
@@
Figure 6: A time lapse of a simple experiment scenario. In (a) the robot pose is a centered on the visual costmap. The global search map, frontier points and non-map points are also shown. The target object is behind the robot starting position. The result of the first planning step is also shown. The planner takes advantage of the concentration of non-map points behind the robot. Figure (b) demonstrates the moment of detection and localization of the target object (suitcase).